# ChronoFlow: A Data-Driven Model for Gyrochronology

### v1

@@ Abstract
Gyrochronology is a technique for constraining stellar ages using rotation periods, which change over a star’s main sequence lifetime due to magnetic braking. This technique shows promise for main sequence FGKM stars, where other methods are imprecise. However, models have historically struggled to capture the observed rotational dispersion in stellar populations. To properly understand this complexity, we have assembled the largest standardized data catalog of rotators in open clusters to date, consisting of $\approx$ 7,400 stars across 30 open clusters/associations spanning ages of 1.5 Myr to 4 Gyr. We have also developed ChronoFlow: a flexible data-driven model which accurately captures observed rotational dispersion. We show that ChronoFlow can be used to accurately forward model rotational evolution, and to infer both cluster and individual stellar ages. We recover cluster ages with a statistical uncertainty of 0.06 dex ($\approx$ 15%), and individual stellar ages with a statistical uncertainty of 0.7 dex. Additionally, we conducted robust systematic tests to analyze the impact of extinction models, cluster membership, and calibration ages on our model’s performance. These contribute an additional  0.06 dex of uncertainty in cluster age estimates, resulting in a total error budget of 0.08 dex. We estimate ages for the NGC 6709 open cluster and the Theia 456 stellar stream, and calculate revised rotational ages for M34, NGC 2516, NGC 1750, and NGC 1647. Our results show that ChronoFlow can precisely estimate the ages of coeval stellar populations, and constrain ages for individual stars. Furthermore, its predictions may be used to inform physical spin down models. ChronoFlow will be publicly available at https://github.com/philvanlane/chronoflow.
@@ 1 Introduction
Lithium depletion, which occurs via nuclear reactions in stellar cores, offers yet another way to estimate ages. Lower-mass stars deplete lithium slower because they are cooler, so the Lithium Depletion Boundary (LDB; i.e. the lowest mass stars that have fully consumed their lithium) serves as an age indicator for stellar populations. Analyzing the LDB (e.g., Jeffries 2014; Beck et al. 2017; Deliyannis et al. 2019) has worked well for stars and clusters $\lesssim$ 200 Myr), but requires spectroscopy to measure Lithium equivalent widths ($EW_{Li}$), and is not well constrained at older ages due to the flattening of Lithium depletion. As an example, Jeffries et al. (2023b) used an empirical model to estimate individual stellar ages and cluster ages up to 6 Gyr. They achieved precisions of 0.1 dex in the best-constrained cases (M dwarfs ¡ 100 Myr), but noted that their recovery of training ages struggled at ¿ 1 Gyr.
A stellar dating technique that has proven useful for low mass MS stars is gyrochronology. Gyrochronology is applicable to any stars with a convective envelope ($\sim$ late F and GKM classes, including fully convective M dwarfs), because these stars lose angular momentum over time due to the ejection of material via magnetized winds. This angular momentum loss is observable as a decrease in rotation rate. This theory of “magnetic braking” was first proposed by Schatzman (1962), and refined further by e.g., Kraft (1967), Wilson (1966), and Weber & Davis (1967). Skumanich (1972) proposed the first quantitative empirical model describing this stellar spin down: $P_{rot}\propto\tau^{\frac{1}{2}}$ (where $P_{rot}$ is the rotation period and $\tau$ is the age of a star).
Gyrochronology took another step forward with foundational papers by Barnes (2003, 2007), who proposed a new empirical model featuring “fast” and “slow” evolutionary tracks. In these models, stellar spin down only depended on color (B-V). Barnes (2010) and Barnes & Kim (2010) developed this further into spindown models which are dependent on the Rossby number $R_{o}$ (the ratio of rotation period to convective turnover time) of a star.
@@
    Weakened magnetic braking and change in rotational evolution of field age stars (Angus et al., 2015; van Saders et al., 2016; Hall et al., 2021; Metcalfe et al., 2022; Silva-Beyer et al., 2022).
@@
    Observational scatter in the $P_{rot}$-$\tau$-mass space has not been well constrained by physical or empirical models to date (e.g., Curtis et al. 2019, 2020; Bouma et al. 2023; Lu et al. 2024a).
@@
To achieve this flexibility, we previously presented the first machine learning (ML) based gyrochronology model proof of concept (Van-Lane et al., 2023), and two other recent studies have also probed the flexibility of empirical models. Bouma et al. (2023) developed an interpolation framework which fits the fast and slow rotator sequences to two normal distributions. Lu et al. (2024a) implemented a Gaussian Process calibrated against seven open clusters and gyro-kinematic ages for $\approx$ 46,000 field stars. While these methods capture some of the dispersion seen in observations, they still require constraints on the models’ functional form.
@@
In this work we present our updated model: ChronoFlow, which does not impose any constraints on the functional form of rotational evolution, and accurately captures observational scatter. We have expanded on our data catalog from Van-Lane et al. (2023) by: (i) including literature rotation periods for 30 open clusters and associations; (ii) standardizing to Gaia DR3 photometry; and (iii) de-reddening using 3D dustmaps from Edenhofer et al. (2023) and Green et al. (2019). We have also optimized our training architecture and enhanced the model itself, which now incorporates photometric uncertainties and cluster membership probabilities.
The rest of the paper is structured as follows. §2 outlines how we have assembled our final catalog, crossmatched all rotation periods to Gaia DR3 (Gaia Collaboration et al., 2023), applied photometric uncertainties, and assigned cluster membership probabilities to all stars. §3 describes the statistical framework for our model, and §4 details the model itself. §5 presents our results, including age recovery tests. In §6 we describe all of the systematic uncertainties that we have characterized, and quantify our model’s performance. We outline applications for ChronoFlow in §7, including new age estimates for five clusters and one stellar stream, and we conclude in §8.
@@ 2 Catalog of Open Cluster Rotators
    Rotation periods. We compiled rotation periods ($P_{rot}$) from 15 literature catalogs. Many of these are themselves compilations, so we trace all rotation periods back to their original measurements. Details are discussed in §2.1.
@@
    Cluster ages. We conducted a literature review of age estimates for the clusters in our catalog; these are compiled in Table 2 and Figure 4.
@@ 2.1 Rotation periods
We started with 15,723 rotation periods compiled from all sources in Table LABEL:tab:prot_source_table. After trimming this list based on quality cuts recommended in the source papers and single-star criteria from each data catalog, 11,427 rotation periods remained. The details of these cuts are described in Appendix B. Since many of our clusters are in multiple data catalogs, there were two types of measurement duplication in these 11,427 periods:
@@
To handle the first case, we identified the source of rotation measurement for all stars across the catalogs. In 420 cases, the same $P_{rot}$ measurement was cited in multiple catalogs (see Appendix C), meaning that 11,007 out of our 11,427 rotation periods were unique measurements.
For the second case, we used the results of our Gaia DR3 crossmatch (see §2.2) to identify stars with rotation periods measured in multiple studies. We found that of the 11,007 measurements, there were 8,249 unique DR3 sources. Of these, 6,069 had a single $P_{rot}$ measurement, 1,403 stars had two measurements, and 777 stars had three measurements. We include all of the stars with a single $P_{rot}$ in our catalog. For the stars with two $P_{rot}$ measurements, we drop those having a fractional difference ($d_{f}$) of $\geq$ 20% from our catalog (68 stars). Figure 1 visualizes this cut. While there was evidence that some of these might have been harmonics (where one $P_{rot}$ is twice or half the value of the other), most did not appear to be. For stars having a fractional difference ¡ 20%, we took the average of the two as our $P_{rot}$.
For stars with three $P_{rot}$ measurements, we first computed the fractional standard deviation of the three values. If it was $\leq$ 2%, we used the average. If not, we compared the difference between the maximum value and middle value ($d_{max}$) to the difference between the minimum value and the middle value ($d_{min}$). If $d_{max}>3d_{min}$ or vice versa, we discarded the outlier. We then considered the fractional difference between the two remaining values. If this was $\geq$ 20%, we exclude the star from our catalog. If it was ¡ 20%, we used the average of the two as our $P_{rot}$. If $d_{max}$ and $d_{min}$ did not differ by more than 3×3$\times$3 × in either direction, we excluded the star if the fractional standard deviation of the three values was $\geq$ 20%, and used the average $P_{rot}$ otherwise.
@@
Total 15 Data Catalogs 36 $P_{rot}$ Sources 20 Source Surveys 11,427 Periods
@@
With this process, only one star with three measured rotation periods was excluded from our catalog, and we were left with 8,181 stars with reliable rotation periods in total. We describe edge cases that we encountered when compiling these rotation periods in Appendix C.
Figure 2: De-reddened color-rotation space for all stars in our final catalog. Stars are color-coded according to the age of their cluster. Older stars appear to have generally converged onto slowly rotating sequences more than younger stars.
@@ 2.2 Photometry
For consistency, we use Gaia DR3 BP and RP bands (Riello et al., 2021) for all photometry. Of the 11,427 compiled rotation periods, we could not find a DR3 crossmatch for 49 of them. Additionally, 170 of the stars that we did crossmatch belong in Taurus, Upper Scorpius, or $\rho$ Ophiucus, but were not part of the Ratzenböck et al. (2023) or Krolikowski et al. (2021) catalogs (which we use as our source of truth for those regions). We successfully crossmatched the other 11,208 rotators, using the steps described in Appendix D, corresponding to 8,181 unique DR3 sources. In our final training catalog we have excluded 21 sources that do not have both BP and RP photometry, and an additional 713 having Gaia RUWE values $\geq$ 1.4 (although these are included in the online table). This last criteria is intended to exclude likely binaries that made it through the source catalog cuts, although we recognize that there are caveats. This cut may not capture short-period binaries (e.g., El-Badry 2024; Grondin et al. 2024), and Penoyre et al. (2022) indicate that a cut of 1.25 is more appropriate for DR3. We include an outlier component in our model to account for binaries that remain in our sample, but acknowledge that handling this more thoughtfully in the future could help our model performance (see §4).
Finally, we apply a cut to exclude non-MS stars. While many of the source catalogs have already been vetted for these, we have performed a final CMD check, and excluded an additional 37 stars (see Appendix A). After these cuts, our final catalog consists of 7,407 unique Gaia DR3 MS sources with reliable rotation periods. These are plotted in Figure 2.
Figure 3: Distribution of $(BP-RP)_{0}$ uncertainties. There is a minimum value of 0.01415 indicated by the red dashed line; this is the DR3 systematic floor corresponding to errors of 0.01 in the BP and RP bands. This plot is truncated at 0.2 mag, but there is a very thin tail that extends to 1.24 mag.
@@ 2.2.1 Stellar de-reddening
We used the Gaia DR3 R.A. and Decl. values to query the dust maps using the dustmaps [^3] https://github.com/gregreen/dustmaps package from Green (2018). We sampled 10 distances for each star from a normal distribution in parallax, after applying zeropoint corrections using gaiadr3_zeropoint (Lindegren et al. 2021). For DR3 sources having a relative parallax error of ¿20%, we instead sampled from a normal distribution in distance representative of the stars in the same cluster having ¡20% fractional parallax error. We also sampled extinctions from each dustmap, and propagated these uncertainties in distance and extinction through the conversions in Eqns. (1) and (2) to get our de-reddened uncertainties. Figure 3 shows the final distribution of these $(BP-RP)_{0}$ uncertainties, truncated at 0.2 mag for ease of visualization.
@@ 2.4 Cluster ages
Figure 4: Literature age estimates for our clusters, demonstrating the coverage in our catalog. The values from the four main catalogs are highlighted explicitly as red, green, pink, and orange points, and the average value is shown as a black diamond (with an error bar showing the standard deviation across the four catalogs). Other literature estimates are shown as translucent blue circles (see Figure 5 as an example). We exclude the subgroups of Taurus and Upper Scorpius/$\rho$ Ophiucus from this plot.
@@
We also conducted a literature review of other age estimates for our clusters, which are summarized in Figure 4 and Table 2. As an example of the variation among studies, Figure 5 presents all age estimates for .
@@
Methodology The method used to estimate the age of the cluster.
Details Family of isochrones used to estimate the age (where applicable).
Age (Myr) The cluster’s age in millions of years.

### v2

@@ Abstract
Gyrochronology is a technique for constraining stellar ages using rotation periods, which change over a star’s main sequence lifetime due to magnetic braking. This technique shows promise for main sequence FGKM stars, where other methods are imprecise. However, the observed dispersion in rotation rates for similar coeval stars has historically been difficult to characterize. To properly understand this complexity, we have assembled the largest standardized data catalog of rotators in open clusters to date, consisting of $\approx$ 7,600 stars across 30 open clusters/associations spanning ages of 1.5 Myr to 4 Gyr. We have also developed ChronoFlow: a flexible data-driven model which accurately captures observed rotational dispersion. We show that ChronoFlow can be used to accurately forward model rotational evolution, and to infer both cluster and individual stellar ages. We recover cluster ages with a statistical uncertainty of 0.06 dex ($\approx$ 15%), and individual stellar ages with a statistical uncertainty of 0.7 dex. Additionally, we conducted robust systematic tests to analyze the impact of extinction models, cluster membership, and calibration ages. These contribute an additional  0.06 dex of uncertainty in cluster age estimates, resulting in a total error budget of 0.08 dex. We apply ChronoFlow to estimate ages for M34, NGC 2516, NGC 6709, and the Theia 456 stellar stream. Our results show that ChronoFlow can precisely estimate the ages of coeval stellar populations, and constrain ages for individual stars. Furthermore, its predictions may be used to inform physical spin down models. ChronoFlow is publicly available at https://github.com/philvanlane/chronoflow.
@@ 1 Introduction
Lithium depletion, which occurs via nuclear reactions in stellar cores, offers yet another way to estimate ages. Lower-mass stars deplete lithium slower because they are cooler, so the Lithium Depletion Boundary (LDB; i.e. the lowest mass stars that have fully consumed their lithium) serves as an age indicator for stellar populations. Analyzing the LDB (e.g., Jeffries 2014; Beck et al. 2017; Deliyannis et al. 2019; Jeffries et al. 2023b) has shown promise for stars and clusters $\lesssim$ 1 Gyr, but requires spectroscopy to measure Lithium equivalent widths ($EW_{Li}$), and is not well constrained at older ages due to the flattening of Lithium depletion. Jeffries et al. (2023b) used an empirical model to estimate individual stellar ages and cluster ages up to 6 Gyr, and achieved precisions of 0.1 dex in the best-constrained cases (M dwarfs $<$ 100 Myr) but struggled to recover training ages at $>$ 1 Gyr.
A stellar dating technique that has proven useful for low mass MS stars is gyrochronology. Gyrochronology is applicable to any stars with a convective envelope ($\sim$ late F and GKM classes, including fully convective M dwarfs), because these stars lose angular momentum over time due to the ejection of material via magnetized winds. This angular momentum loss is observable as a decrease in rotation rate. This theory of “magnetic braking” was first proposed by Schatzman (1962), and refined further by e.g., Kraft (1967), Wilson (1966), and Weber & Davis (1967). Skumanich (1972) proposed a quantitative empirical model describing this stellar spin down: $P_{rot}\propto\tau^{\frac{1}{2}}$ (where $P_{rot}$ is the rotation period and $\tau$ is the age of a star).
Gyrochronology took another step forward with foundational papers by Barnes (2003, 2007), who proposed a new empirical model featuring “fast” and “slow” evolutionary tracks. In these models, the stellar spin down (change in $P_{rot}$) depends on the color (B-V) and age of a star, as well as scaling factors. Barnes (2010) and Barnes & Kim (2010) developed this further into spindown models dependent on the Rossby number $R_{o}$ (the ratio of rotation period to convective turnover time) of a star.
@@
    Weakened magnetic braking (predominantly in K dwarfs), and a change in rotational evolution of field age stars (Angus et al., 2015; van Saders et al., 2016; Curtis et al., 2019; Hall et al., 2021; Metcalfe et al., 2022; Silva-Beyer et al., 2022).
@@
    Observational scatter in $P_{rot}$ for stars of similar ages and masses/temperatures/colors, particularly at young ages, has been difficult to reproduce with physical or empirical models so far (e.g., Curtis et al. 2019, 2020; Bouma et al. 2023; Lu et al. 2024a).
@@
To achieve this flexibility, we previously presented the first machine learning (ML) based gyrochronology model proof of concept (Van-Lane et al., 2023), and two other recent studies have also probed the flexibility of empirical models. Bouma et al. (2023) developed an interpolation framework which fits the slow rotator sequence to a 7th order polynomial with Gaussian scatter, and the fast rotator sequence as a “residual” (compared to the slow sequence mean model) combination of normal and uniform distributions. Lu et al. (2024a) implemented a Gaussian Process gyrochronology model, calibrated against both open clusters and gyro-kinematic field star ages (which are derived using an age-velocity-dispersion relation for groups of stars with similar absolute magnitudes, rotation periods, Rossby numbers and effective temperatures). While these methods capture some of the dispersion seen in observations, they still require constraints on the models’ functional form.
@@
In this work we present our updated model: ChronoFlow, which models rotational evolution flexibly with a neural network, and accurately captures observational scatter. We have expanded on our data catalog from Van-Lane et al. (2023) by: (i) including literature rotation periods for 30 open clusters and associations; (ii) standardizing to Gaia DR3 photometry; and (iii) de-reddening using 3D dustmaps from Edenhofer et al. (2023) and Green et al. (2019). We have also optimized our training architecture and enhanced the model itself, which now incorporates photometric uncertainties and cluster membership probabilities.
The rest of the paper is structured as follows. §2 outlines how we have assembled our final catalog, crossmatched all rotation periods to Gaia DR3 (Gaia Collaboration et al., 2023), applied photometric uncertainties, and assigned cluster membership probabilities to all stars. §3 describes the statistical framework for our model, and §4 details the model itself. §5 presents our results, including age recovery tests. In §6 we describe all of the systematic uncertainties that we have characterized, and quantify our model’s performance. We outline applications for ChronoFlow in §7, including age estimates for five clusters and one stellar stream, and we conclude in §8.
@@ 2 Catalog of Open Cluster Rotators
    Rotation periods. We compiled rotation periods ($P_{rot}$) from 16 literature catalogs. Many of these are themselves compilations, so we trace all rotation periods back to their original measurements. Details are discussed in §2.1.
@@
    Cluster ages. We conducted a literature review of age estimates for the clusters in our catalog; these are compiled in Table 2.
@@ 2.1 Rotation periods
We started with 16,340 rotation periods compiled from all sources in Table LABEL:tab:prot_source_table. After trimming this list based on quality cuts recommended in the source papers and single-star criteria from each data catalog, 12,039 rotation periods remained. The details of these cuts are described in Appendix B. Since many of our clusters are in multiple data catalogs, there were two types of measurement duplication in these 12,039 periods:
@@
To handle the first case, we identified the source of rotation measurement for all stars across the catalogs. In 666 cases, the same $P_{rot}$ measurement was cited in multiple catalogs (see Appendix C), and 11,310 out of our 12,039 rotation periods were unique measurements.
For the second case, we used the results of our Gaia DR3 crossmatch (see §2.2) to identify stars with rotation periods measured in multiple studies. We found that of the 11,310 measurements, there were 8,490 unique DR3 sources. Of these, 6,348 had a single $P_{rot}$ measurement, 1,685 stars had two measurements, and 457 stars had three measurements. We include all of the stars with a single $P_{rot}$ in our catalog. For the stars with two $P_{rot}$ measurements, we drop those having a fractional difference ($d_{f}$) of $\geq$ 20% from our catalog (77 stars). Figure 1 visualizes this cut. While there was evidence that some of these might have been harmonics (where one $P_{rot}$ is twice or half the value of the other), most did not appear to be. For stars having a fractional difference $<$ 20%, we took the average of the two as our $P_{rot}$.
For stars with three $P_{rot}$ measurements, we first computed the fractional standard deviation of the three values. If it was $\leq$ 2%, we used the average. If not, we compared the difference between the maximum value and middle value ($d_{max}$) to the difference between the minimum value and the middle value ($d_{min}$). If $d_{max}>3d_{min}$ or vice versa, we discarded the outlier. We then considered the fractional difference between the two remaining values. If this was $\geq$ 20%, we exclude the star from our catalog. If it was $<$ 20%, we used the average of the two as our $P_{rot}$. If $d_{max}$ and $d_{min}$ did not differ by more than 3×3$\times$3 × in either direction, we excluded the star if the fractional standard deviation of the three values was $\geq$ 20%, and used the average $P_{rot}$ otherwise.
@@
Fritzewski et al. (2020) Irwin et al. (2008) Blanco telescope at CTIO 309
source catalog Yale telescope at CTIO 303
NGC 2548 Barnes et al. (2015) source catalog AIP STELLA I, IAC 59
@@
Total 16 Data Catalogs 37 $P_{rot}$ Sources 20 Source Surveys 12,039 Periods
@@
With this process, only one star with three measured rotation periods was excluded from our catalog, and we were left with 8,412 stars with reliable rotation periods in total. We describe edge cases that we encountered when compiling these rotation periods in Appendix C.
Figure 2: De-reddened color-rotation space for all stars in our final catalog, including the young subgroups of Taurus and Sco-Cen. Stars are color-coded according to the age of their cluster. Older stars appear to have generally converged onto slowly rotating sequences more than younger stars.
@@ 2.2 Photometry
For consistency, we use Gaia DR3 BP and RP bands (Riello et al., 2021) for all photometry. Of the 12,039 compiled rotation periods, we could not find a DR3 crossmatch for 49 of them. Additionally, 170 of the stars that we did crossmatch belong in Taurus, Upper Scorpius, or $\rho$ Ophiucus, but were not part of the Ratzenböck et al. (2023) or Krolikowski et al. (2021) catalogs (which we use as our source of truth for those regions). We successfully crossmatched the other 11,820 rotators, using the steps described in Appendix D, corresponding to 8,412 unique DR3 sources. In our final training catalog we have excluded 21 sources that do not have both BP and RP photometry, three sources from NGC 3532 for which we were unable to calculate de-reddened colors, and an additional 736 having Gaia RUWE values $\geq$ 1.4 (although these are included in the online table). This last criteria is intended to exclude likely binaries that made it through the source catalog cuts, although we recognize that there are caveats. This cut may not capture short-period binaries (e.g., El-Badry 2024; Grondin et al. 2024), and Penoyre et al. (2022) indicate that a cut of 1.25 is more appropriate for DR3. We include an outlier component in our model to account for binaries that remain in our sample, but acknowledge that handling this more thoughtfully in the future could help our model performance (see §6.1).
Finally, we apply a cut to exclude post-MS stars. While many of the source catalogs have already been vetted for these, we have performed a final CMD check, and excluded an additional 37 stars (see Appendix A). After these cuts, our final catalog consists of 7,615 unique Gaia DR3 MS sources with reliable rotation periods. These are plotted in Figure 2. We do not include a cut to remove pre-MS (PMS) stars. Although the rotational behavior of such stars is different than MS stars “spinning down”, we keep PMS stars in our catalog since they can persist in clusters up to hundreds of Myr in age, and therefore are an important factor in the observed rotational distribution of young clusters.
Figure 3: Distribution of $(BP-RP)_{0}$ uncertainties, including propagation of parallax, dustmap, and conversion errors. There is a minimum value of 0.01415 indicated by the red dashed line; this is the DR3 systematic floor corresponding to errors of 0.01 in the BP and RP bands. This plot is truncated at 0.2 mag, but there is a thin tail that extends to 1.24 mag.
@@ 2.2.1 Stellar de-reddening
We used the Gaia DR3 R.A. and Decl. values to query the dust maps using the dustmaps [^3] https://github.com/gregreen/dustmaps package from Green (2018). We sampled 10 distances for each star from a normal distribution in parallax, after applying zeropoint corrections using gaiadr3_zeropoint (Lindegren et al. 2021). For DR3 sources having a relative parallax error of $>$20%, we instead sampled from a normal distribution in distance representative of the stars in the same cluster having $<$20% fractional parallax error. We also sampled extinctions from each dustmap, and propagated these uncertainties in distance and extinction through the conversions in Eqns. (1) and (2) to get our de-reddened uncertainties. Figure 3 shows the final distribution of these $(BP-RP)_{0}$ uncertainties, truncated at 0.2 mag for ease of visualization.
@@ 2.4 Cluster ages
Figure 4: Literature age estimates for our clusters, demonstrating the coverage in our catalog. The values from the four main catalogs are highlighted explicitly as red, green, pink, and orange points, and the average value is shown as a black diamond (with an error bar showing the standard deviation across the four catalogs). Other literature estimates are shown as translucent blue circles (see Figure 5 as an example). We exclude the young subgroups of Taurus and Upper Scorpius/$\rho$ Ophiucus from this plot.
@@
We also conducted a literature review of other age estimates for our clusters, which are summarized in Figure 4 and Table 2. As an example of the variation among studies, Figure 5 presents all age estimates for (some ages are jittered vertically for visualization purposes).
@@
Methodology The technique used to estimate the age of the cluster.
Age The cluster’s age in millions of years.
Age_err1 Upper uncertainty on the age in millions of years (when available).
Age_err2 Lower uncertainty on the age in millions of years (when available).